The sum of the measures of the angles of any triangle is 180 degrees.In triangle ABC,angles A and B have the same measure ,while the measure of angle C is 90 times larger than each of A and B.What are the measures of angle A,B,C.
Answer by rwm(914)

2A +90A=180
92A=180
a=45/23
b=45/23
c=90*45/23